OpenStudy (anonymous):

Solve In 2 + In x=5. Round to the nearest thousandth, if necessary. Can you show the steps so I can go off this one for my other problems like this

Nnesha (nnesha):

ln rules product rule \[\huge\rm \ln x + \ln y = \ln (x \times y)\] if there is plus sign you can change that to multiplication

OpenStudy (anonymous):

So I just multiply?

Nnesha (nnesha):

so how would you change this \[\huge\rm ln 2 + \ln x = ??\] can be written as ??

OpenStudy (anonymous):

In(2 x x) ??

Nnesha (nnesha):

yep right \[\huge\rm ln (2x) = 5\] now you have to take "e" both side to cancel out ln for example

Nnesha (nnesha):

\[\huge\rm e^{\ln +1} = 1\] e can cancel out ln

OpenStudy (anonymous):

Oh goodness.. Im getting confused, if you did that would the In just cancel out? or

Nnesha (nnesha):

yes right left side ln should cancel out

OpenStudy (anonymous):

okay Im following now

Nnesha (nnesha):

yep right \[\huge\rm e^{ln (2x)} = e^5\] now left side e^{ln} can cancel out

OpenStudy (anonymous):

alright, then do I replace the In with e^5?

Nnesha (nnesha):

nope that stays same

OpenStudy (anonymous):

oh okay

Nnesha (nnesha):

when you cancel out e^{ln} then you should have yep right \[\huge\rm 2x = e^5\] now divide both side by 2

OpenStudy (anonymous):

I always get stuck on this step, so x=

Nnesha (nnesha):

just divide both side 2 \[\huge\rm \frac{ 2x }{ 2 }=\frac{ e^5 }{ 2 }\]

OpenStudy (anonymous):

Thats what I did but I wasnt sure if it was right

Nnesha (nnesha):

it says "round to the nearest thousandth" so use calculator put e^5/2

OpenStudy (anonymous):

74.21?

Nnesha (nnesha):

thousandth

Nnesha (nnesha):

\[\huge\rm \frac{ e^5 }{ 2 } =74.2065795513\] now round to the nearest thousandth

Nnesha (nnesha):

+3 decimal places

OpenStudy (anonymous):

74.206

Nnesha (nnesha):

 5 is nexxt to the zero so 6 suppose to be